Prove the following

If a is odd and b is even, then a2 – b2 is an odd number
1
Expert's answer
2019-08-11T08:10:06-0400


a2- b2 = (a - b)(a+b)

a is odd, so a = 2k + 1, where k belongs to natural numbers

b is even, so b = 2m, where m belongs to natural numbers

So, (a - b)(a + b) = (2k + 1 - 2m)(2k + 1 + 2m) .

2k + 1 - 2m is odd, 2k+1 + 2m is odd.

Odd multiplied by odd is always equal to odd.

So, a2 - b2 = (a - b)(a + b) = (2k + 1 - 2m)(2k + 1 + 2m) is odd.